OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. Answer y - 6 = -3x - 3 y = -3x + 3 y - 6 = -3(x + 1) 3x + y = 3

OpenStudy (lucaz):

you have the slope -3, and the point (-1,6) if you plug them into y-y1=m(x-x1) you get: y-(6)=-3(x-(-1))

OpenStudy (lucaz):

the 1st and 3rd are the same answer